# Trackastra: Transformer-based cell tracking for live-cell microscopy

### v1

@@ Abstract
Cell tracking is an omnipresent image analysis task in live-cell microscopy. It is similar to multiple object tracking (MOT), however, each frame contains hundreds of similar-looking objects that can divide, making it a challenging problem. Current state-of-the-art approaches follow the tracking-by-detection paradigm, i.e. first all cells are detected per frame and successively linked in a second step to form biologically consistent cell tracks. Linking is commonly solved via discrete optimization methods, which require manual tuning of hyperparameters for each dataset and are therefore cumbersome to use in practice. Here we propose Trackastra, a general purpose cell tracking approach that uses a simple transformer architecture to directly learn pairwise associations of cells within a temporal window from annotated data. Importantly, unlike existing transformer-based MOT pipelines, our learning architecture also accounts for dividing objects such as cells and allows for accurate tracking even with simple greedy linking, thus making strides towards removing the requirement for a complex linking step. The proposed architecture operates on the full spatio-temporal context of detections within a time window by avoiding the computational burden of processing dense images. We show that our tracking approach performs on par with or better than highly tuned state-of-the-art cell tracking algorithms for various biological datasets, such as bacteria, cell cultures and fluorescent particles. We provide code at https://github.com/weigertlab/trackastra.
@@ 1 Introduction
Figure 1: Overview of Trackastra. Given frame-by-frame object detections in a live-cell video, object features are extracted from a small temporal window and passed as tokens into an encoder-decoder transformer, to predict pairwise associations $\hat{A}$. We apply a parental softmax normalization on $\hat{A}$ to guide the learning directly towards biologically plausible associations. Finally, we build a candidate graph from $\hat{A}$ and prune it with either a greedy algorithm or discrete optimization to obtain a tracking solution.
@@
While cell tracking has been performed manually during initial experiments of embryo development [44], many specialized semi-automated and fully automated algorithms have been developed over the last two decades [17, 14]. Most current state-of-the-art cell tracking methods follow the tracking-by-detection approach, where first cell instances are detected in each frame of a recorded video, which then in a second step are linked across frames to form biologically valid tracks. In recent years, several robust deep-learning-based methods have emerged for the previously tedious and parameter-sensitive first cell detection step [35, 13, 37, 42, 41]. However, the second linking step is still commonly solved with specialized discrete optimisation methods such as integer linear programming (ILP) that allow to enforce the biological constraints on a graph data structure [18, 36, 2, 15]. The costs for such optimization methods have to be carefully chosen, and can range from simple features such as Euclidean distance or intersection-over-union between objects [17, 46] to explicit motion models [48]. More recently, annotated tracking ground truth videos have been used to learn motion models with convolutional neural networks [29, 15, 10, 11, 12, 25], or to learn detection associations in a local context via graph neural networks [1, 39]. In [2, 47], multiple segmentation hypotheses are considered during the linking optimization problem, and a feasible unique solution is enforced with adequate additional constraints. In [34], the linking step is not formulated as a discrete optimization problem, but instead as a regression of each object’s position in the following frame, which lacks a notion of interactions between tracks.
@@
In this paper, we explore the potential of directly learning association costs between segmented objects using a straightforward transformer architecture. This approach models the all-to-all interactions between objects within a short temporal window, while relying solely on shallow per-object features like position and basic shape features. Importantly, we aim to learn the correct associations even for dividing objects, in order to reduce the reliance on computationally expensive combinatorial optimizers and to potentially allow for a greedy linking post-processing only based on the learned associations. The transformer neural network architecture [50] was originally introduced in the context of language translation, a sequence-to-sequence task, where the input is a set of tokens embedded in a continuous vector space. These are then feed through (self/cross) attention layers that enable the model to reason across all tokens at once. This particular inductive bias has produced state-of-the-art results for various predictive tasks using different elementary tokens, for example image patches as tokens for image classification [7], object detection [3] and segmentation [22], amino acids as tokens for protein structure prediction [19], and keypoints as tokens for image feature matching [24] or tasks on point clouds [26]. Notably, tracking in the context of multiple object tracking (MOT) in the domain of natural images has been recently successfully approached with transformers in [33, 45, 54, 4]. Since the temporal succession of cell detections in a video can be seen as a sequence of tokens, we here investigate whether a purely transformer-based model can be used to learn the correct associations between cell detections in a video, especially in the presence of dividing cells.
@@
Our contributions are as follows: i) We provide Trackastra (Tracking-by-Association with Transformers), which is, to the best of our knowledge, the first plain transformer-based cell tracking method that accounts for dividing objects. Our method is simple and formulates the tracking problem as a direct association prediction task, followed by a greedy linking step, reducing the reliance on the computationally expensive optimizer step whose parameter tuning is a common practical challenge in cell tracking experiments. ii) We show that very simple object features such as position or basic shape features are enough for the subsequent model to learn the correct associations, which removes the need for dedicated visual feature extractors such as pretrained CNNs. iii) To ensure the biological consistency of the learned associations, we introduce a blockwise parental softmax normalization method for the association matrix, which allows for one-to-many associations, but not many-to-one. iv) We evaluate our approach experimentally on three datasets from different modalities and demonstrate that Trackastra even outperforms baselines that are highly tuned for specific model organisms.
@@ 1.1 Related work
Cell tracking in microscopy videos is closely related to the extensively studied multiple object tracking (MOT) problem in computer vision [27], for which in recent years several transformer-based approaches have been proposed [45, 33, 54, 4]. In contrast to MOT, however, the cell tracking problem exhibits two crucial differences that motivate our work [23]: Firstly, raw image frames are typically grayscale and contain thousands of objects with very similar visual appearance that are hard to distinguish locally. Secondly, objects are allowed to split (but generally not to fuse), which introduces a more complex combinatorial solution space and precludes computational efficient tracking approaches such as network flows, as commonly used in MOT [38].
@@ 1.1.1 MOT with transformers
In a MOT transformer tracking-by-detection setting, the elementary units of the linking step are detected objects in individual time frames of a video. In [33, 45, 52, 6] linking is done autoregressively with a transformer, i.e. the attention is computed between all object detection tokens in frame $t$ (the keys) and the aggregated token representations of already linked tracks until frame $t-1$ (the queries). Alternatively, the attention matrix can be computed with bi-directional temporal context between single detections only, considering all detections within a local sliding window [54, 32]. Closest to our work is Global Tracking Transformers [54], which however cannot account for cell divisions and heavily relies on appearance features of each detection, which are extracted with deep convolutional neural networks (CNN) from the raw video frames.
1.1.2 Learning based approaches for cell tracking
In microscopy image analysis, transformer-based tracking has been used in single molecule tracking [53], where most detections look similar like in our problem setting, yet it lacks the support of dividing objects. For dividing objects, explicitly learning associations between objects has been explored with Graph Neural Networks in [39, 1] where object features extracted by a CNN interact via the given structure of a hypothesis graph. This constrains detection or track interactions to a certain locality, in contrast to direct all-to-all interactions in a transformer model.
@@ 2 Method
Our proposed Trackastra method operates on raw image sequences and corresponding detections (or segmentation masks) and uses an encoder-decoder transformer to directly predict the pairwise association matrix $A$ between all detections in a local window of consecutive time frames. Specifically, we construct a token for each object and timepoint within the local window and use the sequence of tokens as input to the transformer. The predicted associations $\hat{A}$ are then used as costs in a candidate track graph that is pruned either greedily or via discrete optimization to obtain the final cell tracks. An overview of the full pipeline is shown in Fig. 1. The following sections describe the dataset and training target construction, the transformer architecture, the loss function, the inference and final link assignment, and implementation details.
@@ 2.1 Dataset and association matrix construction
Let $I_{1},I_{2},\dots,I_{T}\in\operatorname{\mathbb{R}}^{w\times h}$ be an image sequence that is grouped into overlapping windows $S_{1},\dots,S_{T-s+1}\in\operatorname{\mathbb{R}}^{s\times w\times h}$ of size $s$. Each window $S_{k}$ contains a set of detections $\{d_{i}\}$ that each correspond to a time point $t_{i}\in\operatorname{\mathbb{N}}$, a center point $p_{i}\in\operatorname{\mathbb{R}}^{2}$, a segmentation mask $m_{i}\in\{0,1\}^{w\times h}$, and other potential object features $z_{i}\in\operatorname{\mathbb{R}}^{n}$ such as basic shape descriptors or mean image intensity of the instance. The goal of the model is to predict an association probability matrix $\hat{A}=(\hat{a}_{ij})$ between all $d_{i}$ in the window $S_{k}$. To construct the target association matrix $A=(a_{ij})$ the set of detections $\{d_{i}\}$ is matched to the set of ground truth objects $V=\{v_{k}\}$ and their ground truth associations. Each ground truth object again corresponds to a time point $t_{k}$, center point $p_{k}$, and a segmentation mask $m_{k}$ and the tracking associations can be described as a directed tree $G=(V,E)$. An edge $e_{kl}$, $k,l\in V$ exists only if $t_{k}+1=t_{l}$ and the objects $v_{k}$ and $v_{l}$ represent the same cell at different time points, or if $v_{k}$ is the mother cell of $v_{l}$. As a simple matching criterion between detections $d_{i}$ and ground truth objects $v_{k}$ we use
@@ 2.2 Transformer architecture
$\mathcal{A}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}+M\right)V$ (3)
where $Q$ are the projected attention queries, $K$ the projected keys, $V$ the projected values, and $M$ is a mask disabling attention for all token pairs whose distance is larger than a user defined threshold $d_{max}$, i.e. $M_{ij}=0$ if $||p_{i}-p_{j}||_{2}\leq d_{max}$ and $M_{ij}=-\infty$ otherwise. In every attention layer, we additionally use rotary positional embeddings (RoPE [43]) to directly and efficiently inject relative positional information. The encoder $f$ transforms the input tokens using $L$ self-attention layers $\mathcal{A}_{f}^{\ell}(X,X,X)$ to obtain representations $Y=f(X)$. The decoder $g$ uses $L$ cross-attention layers $\mathcal{A}_{g}^{\ell}(X,Y,Y)$ to obtain a second set of representations $Z=g(X,Y)$. In between attention layers we use a simple two-layer MLP with GeLU activation, layer normalization and add residual connections following [50]. Finally, we apply a two-layer MLP to $Y$ and $Z$ and compute the logits of the association matrix as their outer product
$\hat{A}=\mathrm{MLP}_{Y}(Y)\cdot\mathrm{MLP}_{Z}(Z)^{T}\quad.$ (4)
@@ 2.3 Parental softmax
where $\mathcal{L}_{BCE}$ is the usual element-wise binary cross-entropy loss, $\lambda\in\operatorname{\mathbb{R}}$ is a small fixed parameter (we use $\lambda=10^{-2}$ throughout), and $W$ is a weighting factor for each matrix element. The elementwise weighting terms $W$ are set to
@@ 2.4 Inference and linking
$\bar{a}_{i^{\prime}j^{\prime}}=\frac{1}{s-1}\sum_{i^{\prime},j^{\prime}\in S_{t}}\hat{a}_{i^{\prime}j^{\prime}}^{(S_{t})}\quad.$ (8)
@@ 2.4.1 Greedy
$\mathrm{deg}^{+}(v)\leq 2\qquad\forall v\in V_{s}$
$\mathrm{deg}^{-}(v)\leq 1\qquad\forall v\in V_{s}\quad.$ (9)
@@ 2.5 Implementation details
Training details:
Trackastra can be trained on a single GPU (e.g. Nvidia RTX 4090) for prototypical 2d cell tracking datasets. We set window size $s=6$, embedding dimension $d=256$, number of encoder and decoder attention layers $L=6$, the maximum number of tokens per window $|D|=2048$, and batch size 8.
Shallow object features:
As basic object features we use the mean intensity, the object area and the inertia tensor of the object region [16].
Augmentations:
We apply the following data augmentations jointly to all frames in a window: flips, shifts, rotations, shear, scaling, and intensity shifting and scaling. The augmentations are applied directly to the object features.
@@ 3 Experiments
In the following we present tracking experiments on three different datasets, each with different characteristics.
@@ 3.1 Metrics
    $EC$: Change the semantics of a predicted edge (linear chain links vs. division links).
@@
$\mathrm{AOGM}=5\cdot|NS|+10\cdot|FN|+1\cdot|FP|+1\cdot|ED|+1.5\cdot|EA|+1\cdot|EC|\quad,$ (10)
which represents the quality of a cell tracking solution for a given video in terms of errors to tolerate, in contrast to the usual dominant fraction of simple correctly assigned links. Note that cell divisions are counted implicitly in AOGM.
We also report the normalized version of AOGM commonly used in cell tracking [49]
@@
where $\mathrm{AOGM_{0}}$ is the AOGM of an empty predicted graph.
Finally, when using predicted segmentations, we also report AOGM$+$, which removes the constant penalty incurred by missing detections.
Figure 3: Error trees on a challenging Bacteria test video. Time on the vertical axis, true positive edges in green, false positive edges in magenta, false negative edges in cyan.

### v2

@@ Abstract
Cell tracking is a ubiquitous image analysis task in live-cell microscopy. Unlike multiple object tracking (MOT) for natural images, cell tracking typically involves hundreds of similar-looking objects that can divide in each frame, making it a particularly challenging problem. Current state-of-the-art approaches follow the tracking-by-detection paradigm, i.e. first all cells are detected per frame and successively linked in a second step to form biologically consistent cell tracks. Linking is commonly solved via discrete optimization methods, which require manual tuning of hyperparameters for each dataset and are therefore cumbersome to use in practice. Here we propose Trackastra, a general purpose cell tracking approach that uses a simple transformer architecture to directly learn pairwise associations of cells within a temporal window from annotated data. Importantly, unlike existing transformer-based MOT pipelines, our learning architecture also accounts for dividing objects such as cells and allows for accurate tracking even with simple greedy linking, thus making strides towards removing the requirement for a complex linking step. The proposed architecture operates on the full spatio-temporal context of detections within a time window by avoiding the computational burden of processing dense images. We show that our tracking approach performs on par with or better than highly tuned state-of-the-art cell tracking algorithms for various biological datasets, such as bacteria, cell cultures and fluorescent particles. We provide code at https://github.com/weigertlab/trackastra.
@@ 1 Introduction
Figure 1: Overview of Trackastra. Given frame-by-frame object detections in a live-cell video, object features are extracted from a small temporal window and passed as tokens into an encoder-decoder transformer, to predict pairwise associations $\hat{A}$. We apply a parental softmax normalization on $\hat{A}$ to guide the learning directly towards biologically plausible associations. Finally, we build a candidate graph by averaging the predictions $\hat{A}$ over a sliding window, and obtain a tracking solution by pruning the graph with either a greedy algorithm or discrete optimization.
@@
While cell tracking has been performed manually during initial experiments of embryo development [48], many specialized semi-automated and fully automated algorithms have been developed over the last two decades [16]. Most current state-of-the-art cell tracking methods follow the tracking-by-detection approach, where first cell instances are detected in each frame of a recorded video, which then in a second step are linked across frames to form biologically valid tracks. In recent years, several robust deep-learning-based methods have emerged for the previously tedious and parameter-sensitive first cell detection step [38, 15, 40, 45, 44]. However, the second linking step is still commonly solved with specialized discrete optimization methods such as integer linear programming (ILP) that allow to enforce the biological constraints on a graph data structure [20, 39, 2, 17]. The costs for such optimization methods have to be carefully chosen, and can range from simple features such as Euclidean distance or intersection-over-union between objects [19, 50] to explicit motion models [52]. More recently, annotated tracking ground truth videos have been used to learn cell flow models with convolutional neural networks for better cost prediction [31, 17, 12, 13, 14, 27, 47], or to directly learn object associations in a local context via graph neural networks [1, 42] or via iterative semantic segmentation [37]. Furthermore in [2, 51], multiple segmentation hypotheses are considered during the discrete linking optimization problem, and a feasible unique solution is enforced with adequate additional constraints.
In this paper, we explore the potential of directly learning association costs between segmented objects using a straightforward transformer architecture. This approach models the all-to-all interactions between objects within a short temporal window, while relying solely on shallow per-object features like position and basic shape features. Importantly, we aim to learn the correct associations even for dividing objects, in order to reduce the reliance on computationally expensive combinatorial optimizers and to potentially allow for a greedy linking post-processing only based on the learned associations.
@@
The transformer neural network architecture [54] was originally introduced in the context of language translation, a sequence-to-sequence task, where the input is a set of tokens embedded in a continuous vector space. These are then fed through (self/cross) attention layers that enable the model to reason across all tokens at once. This particular inductive bias has produced state-of-the-art results for various predictive tasks using different elementary tokens, for example image patches as tokens for image classification [7], object detection [3] and segmentation [24], amino acids as tokens for protein structure prediction [21], and keypoints as tokens for image feature matching [26] or tasks on point clouds [28]. Notably, multiple object tracking (MOT) in the domain of natural images has recently been successfully approached with transformers in [36, 49, 59, 4]. Since the temporal succession of cell detections in a video can be seen as a sequence of tokens, we here investigate whether a purely transformer-based model can be used to learn the correct associations between cell detections in a video, especially in the presence of dividing cells.
@@
Our contributions are as follows: i) We provide Trackastra (Tracking-by-Association with Transformers), which is, to the best of our knowledge, the first plain transformer-based cell tracking method that accounts for dividing objects. Our method is simple, formulating tracking as a direct association prediction task followed by greedy linking, thereby reducing the reliance on the computationally expensive optimizer step whose hyperparameter tuning is a common practical challenge in cell tracking experiments. ii) We show that very simple object features such as position or basic shape features are enough for the subsequent model to learn the correct associations, which removes the need for dedicated visual feature extractors such as pretrained CNNs. iii) To ensure the biological consistency of the learned associations, we introduce a blockwise parental softmax normalization method for the association matrix, which allows for one-to-many associations, but not many-to-one. iv) We evaluate our approach experimentally on three datasets from different modalities and demonstrate that Trackastra outperforms even baselines that are highly tuned for specific model organisms.
@@ 1.1 Related work
Cell tracking in microscopy videos is closely related to the extensively studied multiple object tracking (MOT) problem in computer vision [29], for which in recent years several transformer-based approaches have been proposed [49, 36, 59, 4]. In contrast to MOT [25], however, the cell tracking problem exhibits two crucial differences that motivate our work: First, raw image frames are typically grayscale and contain thousands of objects with very similar visual appearance that are hard to distinguish locally. Second, objects are allowed to split (but generally not to fuse), which introduces a more complex combinatorial solution space and precludes computationally efficient tracking approaches commonly used in MOT, such as network flows [41].
@@ 1.1.1 MOT with transformers
In [36, 49, 57, 6] MOT is done autoregressively end-to-end with a transformer, i.e. the attention is computed between all candidate object detection tokens in frame $t$ (the keys) and the aggregated token representations of already linked tracks until frame $t-1$ (the queries). Alternatively, the attention matrix can be computed with bi-directional temporal context between single detections only, considering all detections within a local sliding window [59, 35]. Closest to our work is the tracking-by-detection approach Global Tracking Transformers [59], which, however, cannot account for cell divisions and heavily relies on appearance features of each detection that are extracted with a convolutional neural network (CNN) from the raw video frames.
1.1.2 Learning based association prediction for cell tracking
In microscopy image analysis, transformer-based tracking has been used in the context of single particle tracking [58] where most detections look similar like in our problem setting, however without supporting dividing objects. In [37], the association step involves densely predicting for each object its mask in the subsequent frame, which is then used to generate appropriate linking costs. For dividing objects, explicitly learning associations between objects has been explored with Graph Neural Networks in [42, 1], where object features extracted by a CNN interact via the given structure of a hypothesis graph. This constrains detection interactions to a certain locality, in contrast to direct all-to-all interactions in a transformer.
@@ 2 Method
Our proposed Trackastra method operates on raw image sequences and corresponding detections (or segmentation masks) and uses an encoder-decoder transformer to directly predict the pairwise association matrix $A$ between all detections in a local window of consecutive time frames. Specifically, we construct a token for each object and timepoint within the local window and use this sequence of tokens as input to the transformer. The predicted associations $\hat{A}$ are then used as costs in a candidate track graph that is pruned either greedily or via discrete optimization to obtain the final cell tracks. An overview of the full pipeline is shown in Fig. 1. The following sections describe the dataset and training target construction, the transformer architecture, the loss function, the inference and final link assignment, and implementation details.
@@ 2.1 Dataset and association matrix construction
Let $I_{1},I_{2},\dots,I_{T}\in\operatorname{\mathbb{R}}^{w\times h}$ be an image sequence that is grouped into overlapping windows $S_{1},\dots,S_{T-s+1}\in\operatorname{\mathbb{R}}^{s\times w\times h}$ of size $s$. Each window $S_{n}$ contains a set of detections $\{d_{i}\}$ that each correspond to a time point $t_{i}\in\operatorname{\mathbb{N}}$, a center point $p_{i}\in\operatorname{\mathbb{R}}^{2}$, a segmentation mask $m_{i}\in\{0,1\}^{w\times h}$, and other potential object features $z_{i}\in\operatorname{\mathbb{R}}^{z}$ such as basic shape descriptors or mean image intensity of the instance. The goal of the model is to predict an association probability matrix $\hat{A}=(\hat{a}_{ij})$ between all $d_{i}$ in the window $S_{n}$. To construct the target association matrix $A=(a_{ij})$ the set of detections $\{d_{i}\}$ is matched to the set of ground truth objects $V=\{v_{k}\}$ and their ground truth associations. Each ground truth object again corresponds to a time point $t_{k}$, center point $p_{k}$, and a segmentation mask $m_{k}$ and the tracking associations can be described as a directed tree $G=(V,E)$. An edge $e_{kl}$, $k,l\in V$ exists only if $t_{k}+1=t_{l}$ and the objects $v_{k}$ and $v_{l}$ represent the same cell at different time points, or if $v_{k}$ is the mother cell of $v_{l}$. As a simple matching criterion between detections $d_{i}$ and ground truth objects $v_{k}$ we use
@@ 2.2 Transformer architecture
$\mathcal{A}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}+M\right)V\quad,$ (3)
where $Q$ are the projected attention queries, $K$ the projected keys, $V$ the projected values, and $M$ is a mask disabling attention for all token pairs whose distance is larger than a user defined threshold $d_{max}$, i.e. $M_{ij}=0$ if $||p_{i}-p_{j}||_{2}\leq d_{max}$ and $M_{ij}=-\infty$ otherwise. In every attention layer, we additionally use rotary positional embeddings (RoPE [46]) to directly and efficiently inject relative spatial and temporal information. The encoder $f$ transforms the input tokens using $L$ self-attention layers $\mathcal{A}_{f}^{\ell}(X,X,X)$ to obtain representations $Y=f(X)$. The decoder $g$ uses $L$ cross-attention layers $\mathcal{A}_{g}^{\ell}(X,Y,Y)$ to obtain a second set of representations $Z=g(X,Y)$. In between attention layers we use a simple two-layer MLP with GeLU activation, layer normalization and add residual connections following [54]. Finally, we apply two-layer MLPs to $Y$ and $Z$ and compute the logits $\hat{A}$ of the final association matrix as the outer product $\hat{A}=\mathrm{MLP}_{Y}(Y)\cdot\mathrm{MLP}_{Z}(Z)^{T}$.
@@ 2.3 Parental softmax
where $\mathcal{L}_{BCE}$ is the usual element-wise binary cross-entropy loss, $\lambda\in\operatorname{\mathbb{R}}$ is a small fixed hyperparameter (we use $\lambda=10^{-2}$ throughout), and $W$ is a weighting factor for each matrix element. The elementwise weighting terms $W$ are set to
@@ 2.4 Inference and linking
$\bar{a}_{i^{\prime}j^{\prime}}=\frac{1}{s-1}\sum_{\{S_{n}|i^{\prime},j^{\prime}\in S_{n}\}}\tilde{a}_{i^{\prime}j^{\prime}}^{(S_{n})}\quad.$ (7)
@@ 2.4.1 Greedy
$\mathrm{deg}^{+}(v)\leq 2\quad\forall v\in V_{S}\quad,\qquad\mathrm{deg}^{-}(v)\leq 1\quad\forall v\in V_{S}\>.$ (8)
@@ 2.5 Implementation details
Trackastra can be trained on a single GPU (e.g. Nvidia RTX 4090) for prototypical 2D cell tracking datasets. The transformer is implemented in PyTorch. We set window size $s=6$, embedding dimension $d=256$, number of encoder and decoder attention layers $L=6$, the maximum number of tokens per window $|D|=2048$, and batch size 8. As shallow object features $z_{i}$ we use the mean intensity, the object area and the inertia tensor of the object region [18]. We apply the following data augmentations jointly to all frames in a window: flips, shifts, rotations, shear, scaling, intensity shifting and scaling, and temporal subsampling. The augmentations are applied directly to the object features. Trackastra association prediction scales to videos with thousands of objects, e.g., inference runs at $\sim 1$ FPS on a single GPU for 2k objects per frame. The ILP is implemented in Motile [10].
@@ 3 Experiments
In the following we present tracking experiments on three datasets with varying characteristics, preceded by the employed evaluation measures.
@@ 3.1 Metrics
    $EC$: Change semantics of predicted edge (linear chain link vs. division link).
@@
$\mathrm{AOGM}=5\cdot|NS|+10\cdot|FN|+1\cdot|FP|+1\cdot|ED|+1.5\cdot|EA|+1\cdot|EC|\>,$ (9)
which represents the quality of a cell tracking solution for a given video in terms of errors to tolerate, in contrast to the usual dominant fraction of simple correctly assigned links. Note that cell divisions are counted implicitly in AOGM. We also report $\mathrm{TRA}$, which is the normalized version of AOGM commonly used in cell tracking [53]
@@
where $\mathrm{AOGM_{0}}$ is the AOGM of an empty predicted graph. Finally, when using predicted segmentations, we also report AOGM$+$, which removes the constant penalty incurred by missing detections.
Figure 3: Error trees on a challenging Bacteria test video. Time on the vertical axis, edges colored as true positive (green), false positive (magenta) and false negative (cyan).